# Weighted Point Configurations with Hyperuniformity: An Ecological Example and Models

### v2

@@ 1 Introduction
Recently in condensed matter physics and related material sciences, correlated particle systems are said to be in a hyperuniform state, when density fluctuations are anomalously suppressed in large-scale [10]. Consider a bounded domain with linear size $\ell$ in the $d$-dimensional space, which we call an observation window. In this window we measure the mean number of points $\langle N_{\ell}\rangle$ and its variance $\sigma^{2}_{\ell}$. In a hyperuniform system, as the window size $\ell\to\infty$, $\sigma^{2}_{\ell}$ grows more slowly than $\langle N_{\ell}\rangle$ which is proportional to the window volume. Typical disordered systems (e.g., liquids and structured glasses) have the standard scaling behavior $\sigma_{\ell}\sim\ell^{d}$ as $\ell\to\infty$, and hence they are not hyperuniform. Periodic point configuration (e.g., lattice points of perfect crystals) are obviously hyperuniform systems, since the number fluctuations are concentrated near the window boundary and hence have the surface-area scaling $\sigma^{2}_{\ell}\sim\ell^{d-1}$ as $\ell\to\infty$. We are interested in the hyperuniform point processes, which are not perfect lattice, but are randomly distributed following some probability laws. A variety of examples of disordered or random point processes showing hyperuniformity have been reported in the review paper by Torquato [10]. A typical example on the two-dimensional plane, which has been well studied in mathematical physics, is the complex Ginibre point process (GPP) [4]. This is realized as the statistical ensemble of eigenvalues of non-Hermitian random matrices on the complex plane $\mathbb{C}$, which can be identified with the two-dimensional space $\mathbb{R}^{2}$. There first we consider an $n\times n$ matrix whose entries are all independently and identically distributed (i.i.d.) complex standard Gaussian random variables. We have a nearly uniform distribution of eigenvalues in a disk centered at the origin with radius proportional to $\sqrt{n}$ on $\mathbb{C}$. If we consider the $n\to\infty$ limit, we have a translation-invariant point process on $\mathbb{C}$ [4]. It was proved that the obtained GPP is hyperuniform with $\sigma_{\ell}\sim\ell$ [11]. The origin of such hyperuniformity in a random point configuration is long-ranged repulsive interaction between any pair of points. As a matter of fact, GPP is a determinantal point process (DPP) [8], for which we can show that a logarithmic repulsive potential acts between any pair of points [3, 12]. It should be remarked that DPP is also called fermion point process [13], since the determinantal structure of correlation functions is essentially equivalent with the Slater determinant describing wave function of fermionic particle systems. We notice that GPP has been studied as a typical example of strongly correlated point process in two dimensions in non-Hermitian physics [14], spectral theory [15], random matrix theory [16], and so on, and used in many applications (see, for instance, the references [17, 18, 19, 20]).
@@
Figure 1: A sample of point configuration of center-of-masses of bushes in a desert. The sizes of area are indicated not by real lengths in meters but by numbers of pixels for the digital data; $L_{x}=1627$, $L_{y}=895$. The total number of points is $N=9853$ and the density is $\rho\fallingdotseq 6.77\times 10^{-3}$. The bush-size (mass) distribution is represented by colors of each points; the points with masses $1\leq M(\bm{X}_{j})\leq 49$ are green, the points with $50\leq M(\bm{X}_{j})\leq 99$ are blue, and the points with $M(\bm{X}_{j})\geq 100$ are red. See Section 3 showing how to produce such digital data.
@@ 2.1 Marked and unmarked point processes
since (2) gives a total number of points included in $B$ for any subset $B\subset S$; $\Xi(B)=\sum_{\bm{X}:\bm{X}\in\mathcal{P}\cap B}1$. We call $\Xi$ associated with $\mathcal{P}$ a unmarked point process, or simply a point process. We also consider a marked point process, in which each point $\bm{X}\in\mathcal{P}$ carries a variable $M(\bm{X})$. Such a marked point process is denoted by
@@ 3.1 Bush distributions in deserts
We will call the sample shown by Fig, 1 simply (Argentina) bush 1 in the following. By the similar procedure explained above, we have prepared other four samples from the same desert in Argentina, two samples from a desert in Algeria, one from Australia, and one from Kenya. The detailed data are listed out in Table 1: The areas of Nos.1–5 are in the Talampaya Natural Park, La Rioja, Argentina, Nos.5 and 6 in the Parc culturel national de Tindouf, Oum El Assel, Algeria, No. 7 at Flying Fox, Northrm Territory, Australia, and No. 8 at Maikona, Kenya. The latitude and longitude are for the center of each area. Real sizes in the east-west (EW), and north-south (NS) directions are given in the metric unit, whose pixels in the digital data are also shown. The total number of bushes in each area is given by $N$. The areas of Algeria 1, 2 are found in a Wadi and hence the aspect ratios $\lambda=\Lambda_{y}/\Lambda_{x}$ are quite deviated from 1.
@@ 3.3 Measurements of hyperuniformity by finite-size data
We prepare a series of windows parameterized by linear size $\ell$, $\{\Lambda_{\ell}\}$, such that $\Lambda_{\ell}$ expands monotonically as $\ell$ increases and $\Lambda_{\ell}$ will cover the whole space asymptotically in the limit $\ell\to\infty$. As explained in Section 2.2, the hyperuniformity of random point process is defined and classified mathematically by asymptotices in the limit $\ell\to\infty$ of the ratio $R_{\ell}$ of the variance with respect to the expectation concerning the number of points included $\Lambda_{\ell}$. In real experiments of physics as well as observation in ecological systems, however, we have only a few number of sample configurations in finite regions.
@@
Figure 4: The region $L_{x}\times L_{y}$ on which a sample of point process is given is divided into $m\times m$ subdomains. Each subregion is a similar rectangular $\ell_{x}\times\ell_{y}$ denoted by $\Lambda_{\ell_{x}}^{j}$, $j=1,2,\dots,m^{2}$. The case with $m=5$ is shown.
@@ 3.4 Hyperuniformity of marked point processes of bushes in a desert in Argentina
In Fig. 5 (a), we show dependence on $\log\ell_{x}$ of logarithms of $R^{\rm point}_{\ell_{x}}$ for the unmarked point configurations of the sample bush 1, $\log R^{\rm PPP}_{\ell_{x}}$ and $\log R^{\rm GPP}_{\ell x}$. For the totally five samples from the desert in Argentina, bush 1–5 in Table 1, $\log R^{\rm mass}_{\ell_{x}}$ are also plotted versus $\log\ell_{x}$ in Fig. 5 (a), which are for the marked point processes. In the numerical measurements of variances by (8) and (10) in Section 3.3, the numbers of subregions $m$ become relatively small for large values of $\ell_{x}$, and hence the obtained plots are scattered. For this reason we only use the data with $\log\ell_{x}<4.5$.
@@ 4.2.1 Hyperuniformity
First we explain the results for the marked point processes shown by Fig. 9 (a). For fixed $L=2000$, as $a_{0}$ increases from 500 to 4000, the linear region extends systematically. For the largest value $a_{0}=4000$, the linear region also systematically extends as $L$ increases from 1000 to 4000. By these results, we think that if both of $L$ and $a_{0}$ are sufficiently large, and if the number of iteration-terms $T$ of the transformation $\mathcal{T}$ is large enough, our procedure produced marked point processes with hyperuniformity. The linear fitting to (12) of the dots of $L=4000$ and $a_{0}=4000$ gives

### v3

@@ 1 Introduction
Recently in condensed matter physics and related material sciences, correlated particle systems are said to be in a hyperuniform state, when density fluctuations are anomalously suppressed in large-scale [10]. Consider a bounded domain with linear size $\ell$ in the $d$-dimensional space, which we call an observation window. In this window we measure the mean number of points $\langle N_{\ell}\rangle$ and its variance $\sigma^{2}_{\ell}$. In a hyperuniform system, as the window size $\ell\to\infty$, $\sigma^{2}_{\ell}$ grows more slowly than $\langle N_{\ell}\rangle$ which is proportional to the window volume. Typical disordered systems (e.g., liquids and structured glasses) have the standard scaling behavior $\sigma^{2}_{\ell}\simeq\ell^{d}$ as $\ell\to\infty$, and hence they are not hyperuniform. Periodic point configuration (e.g., lattice points of perfect crystals) are obviously hyperuniform systems, since the number fluctuations are concentrated near the window boundary and hence have the surface-area scaling $\sigma^{2}_{\ell}\simeq\ell^{d-1}$ as $\ell\to\infty$. We are interested in the hyperuniform point processes, which are not perfect lattice, but are randomly distributed following some probability laws. A variety of examples of disordered or random point processes showing hyperuniformity have been reported in the review paper by Torquato [10]. A typical example on the two-dimensional plane, which has been well studied in mathematical physics, is the complex Ginibre point process (GPP) [4]. This is realized as the statistical ensemble of eigenvalues of non-Hermitian random matrices on the complex plane $\mathbb{C}$, which can be identified with the two-dimensional space $\mathbb{R}^{2}$. First we consider an $n\times n$ matrix whose entries are all independently and identically distributed (i.i.d.) complex standard Gaussian random variables. We have a nearly uniform distribution of eigenvalues in a disk centered at the origin with radius proportional to $\sqrt{n}$ on $\mathbb{C}$. If we consider the $n\to\infty$ limit, we have a translation-invariant point process on $\mathbb{C}$ [4]. It was proved that the obtained GPP is hyperuniform with $\sigma^{2}_{\ell}\simeq\ell$ [11]. The origin of such hyperuniformity in a random point configuration is long-ranged repulsive interaction between any pair of points. As a matter of fact, GPP is a determinantal point process (DPP) [8], for which we can show that a logarithmic repulsive potential acts between any pair of points [3, 12]. It should be remarked that DPP is also called fermion point process [13], since the determinantal structure of correlation functions is essentially equivalent with the Slater determinant describing wave function of fermionic particle systems. We notice that GPP has been studied as a typical example of strongly correlated point process in two dimensions in non-Hermitian physics [14], spectral theory [15], random matrix theory [16], and so on, and used in many applications (see, for instance, the references [17, 18, 19, 20]).
@@
Figure 1: A sample of point configuration of center-of-masses of bushes in a desert. The sizes of area are indicated not by real lengths in meters but by numbers of pixels for the digital data; $L_{x}=1626$, $L_{y}=895$. The total number of points is $N=9853$ and the density is $\rho\fallingdotseq 6.77\times 10^{-3}$. The bush-size (mass) distribution is represented by colors of each points; the points with masses $1\leq M(\bm{X}_{j})\leq 49$ are green, the points with $50\leq M(\bm{X}_{j})\leq 99$ are blue, and the points with $M(\bm{X}_{j})\geq 100$ are red. See Section 3 showing how to produce such digital data.
@@ 2.1 Marked and unmarked point processes
since (2) gives a total number of points included in $B$ for any subset $B\subset S$; $\Xi(B)=\sum_{\bm{X}:\bm{X}\in\mathcal{P}\cap B}1$. We call $\Xi$ associated with $\mathcal{P}$ an unmarked point process, or simply a point process. We also consider a marked point process, in which each point $\bm{X}\in\mathcal{P}$ carries a variable $M(\bm{X})$. Such a marked point process is denoted by
@@ 3.1 Bush distributions in deserts
We will call the sample shown by Fig, 1 simply (Argentina) bush 1 in the following. By the similar procedure explained above, we have prepared other four samples from the same desert in Argentina, two samples from a desert in Algeria, one from Australia, and one from Kenya. The detailed data are listed out in Table 1: The areas of Nos.1–5 are in the Talampaya Natural Park, La Rioja, Argentina, Nos.5 and 6 in the Parc culturel national de Tindouf, Oum El Assel, Algeria, No. 7 at Flying Fox, Northrm Territory, Australia, and No. 8 at Maikona, Kenya. The latitude and longitude are for the center of each area. Real sizes in the east-west (EW), and north-south (NS) directions are given in the metric unit, whose pixels in the digital data are also shown. The total number of bushes in each area is given by $N$. The areas of Algeria 1, 2 are found in a Wadi and hence the aspect ratios $\lambda=L_{y}/L_{x}$ are quite deviated from 1.
@@ 3.3 Measurements of hyperuniformity by finite-size data
We prepare a series of windows parameterized by linear size $\ell$, $\{\Lambda_{\ell}\}$, such that $\Lambda_{\ell}$ expands monotonically as $\ell$ increases and $\Lambda_{\ell}$ will cover the whole space asymptotically in the limit $\ell\to\infty$. As explained in Section 2.2, the hyperuniformity of random point process is defined and classified mathematically by asymptotics in the limit $\ell\to\infty$ of the ratio $R_{\ell}$ of the variance with respect to the expectation concerning the number of points included in $\Lambda_{\ell}$. In real experiments of physics as well as observation in ecological systems, however, we have only a few number of sample configurations in finite regions.
@@
Figure 4: The region $L_{x}\times L_{y}$ on which a sample of point process is given is divided into $m\times m$ subregions. Each subregion is a similar rectangular $\ell_{x}\times\ell_{y}$ denoted by $\Lambda_{\ell_{x}}^{j}$, $j=1,2,\dots,m^{2}$. The case with $m=5$ is shown.
@@ 3.4 Hyperuniformity of marked point processes of bushes in a desert in Argentina
In Fig. 5 (a), we show dependence on $\log\ell_{x}$ of logarithms of $R^{\rm point}_{\ell_{x}}$ for the unmarked point configurations of the sample bush 1, $\log R^{\rm PPP}_{\ell_{x}}$ and $\log R^{\rm GPP}_{\ell x}$. For the totally five samples from the desert in Argentina, bush 1–5 in Table 1, $\log R^{\rm mass}_{\ell_{x}}$ are also plotted versus $\log\ell_{x}$ in Fig. 5 (a), which are for the marked point processes. In the numerical measurements of variances by (8) and (10) in Section 3.3, the numbers of subregions $m^{2}$ become relatively small for large values of $\ell_{x}$, and hence the obtained plots are scattered. For this reason we only use the data with $\log\ell_{x}<4.5$.
@@ 4.2.1 Hyperuniformity
First we explain the results for the marked point processes shown by Fig. 9 (a). For fixed $L=2000$, as $a_{0}$ increases from 500 to 4000, the linear region extends systematically. For the largest value $a_{0}=4000$, the linear region also systematically extends as $L$ increases from 1000 to 4000. By these results, we think that if both $L$ and $a_{0}$ are sufficiently large, and if the number of iteration-terms $T$ of the transformation $\mathcal{T}$ is large enough, our procedure produced marked point processes with hyperuniformity. The linear fitting to (12) of the dots of $L=4000$ and $a_{0}=4000$ gives